Kohler Corporation reports the following components of stockholders’ equity on December 31, 2016:

Common stock—$25 par value, 100,000 shares authorized,
40,000 shares issued and outstanding
$ 1,000,000
Paid-in capital in excess of par value, common stock 80,000
Retained earnings 460,000
Total stockholders' equity $ 1,540,000


In year 2017, the following transactions affected its stockholders’ equity accounts.

Jan. 1 Purchased 5,000 shares of its own stock at $20 cash per share.
Jan. 5 Directors declared a $2 per share cash dividend payable on February 28 to the February 5 stockholders of record.
Feb. 28 Paid the dividend declared on January 5.
July 6 Sold 1,875 of its treasury shares at $24 cash per share.
Aug. 22 Sold 3,125 of its treasury shares at $17 cash per share.
Sept. 5 Directors declared a $2 per share cash dividend payable on October 28 to the September 25 stockholders of record.
Oct. 28 Paid the dividend declared on September 5.
Dec. 31 Closed the $388,000 credit balance (from net income) in the Income Summary account to Retained Earnings.


Required:

1. Prepare journal entries to record each of these transactions for 2017.
2. Prepare a statement of retained earnings for the year ended December 31, 2017.
3. Prepare the stockholders' equity section of the company’s balance sheet as of December 31, 2017.
Answer #1
1
Date General Journal Debit Credit
1-Jan Treasury stock, Common 100,000 =5000*20
Cash 100,000
5-Jan Retained earnings 70,000 =(40000-5000)*2
Common dividend payable 70,000
28-Feb Common dividend payable 70,000
Cash 70,000
6-Jul Cash 45,000 =1875*24
Treasury stock, Common 37,500 =1875*20
Paid-in capital, Treasury stock 7,500
22-Aug Cash 53,125 =3125*17
Paid-in capital, Treasury stock 7,500
Retained earnings 1,875
Treasury stock, Common 62,500 =3125*20
5-Sep Retained earnings 80,000 =40000*2
Common dividend payable 80,000
28-Oct Common dividend payable 80,000
Cash 80,000
31-Dec Income summary 388,000
Retained earnings 388,000
2
KOHLER CORPORATION
Statement of Retained Earnings
For Year Ended December 31, 2017
Retained earnings, Dec. 31, 2016 460,000
Add: Net income 388,000
848,000
Less: Cash dividends declared -150,000
Less: Treasury stock reissuances -1,875
Retained earnings, Dec. 31, 2017 696,125
3
KOHLER CORPORATION
Stockholders' Equity Section of the Balance Sheet
December 31, 2017
Common stock - $25 par value 1,000,000
Paid-in capital in excess of par value, common stock 80,000
Total contributed capital 1,080,000
Retained earnings 696,125
Total stockholders' equity 1,776,125